Exercise 21-15 Assume that on January 1, 2017, Elmers Restaurants sells a computer system to Liquidity Finance Co. for $709,000 and immediately leases the computer system back. The relevant information is as follows. 1. The computer was carried on Elmers books at a value of $636,000. 2. The term of the noncancelable lease is 10 years; title will transfer to Elmer 3. The lease agreement requires equal rental payments of $115,386 at the end of each year 4. The incremental borrowing rate for Elmer is 12%. Elmer is aware that Liquidity Finance Co. set the annual rental to insure a rate of return of 10%. 5. The computer has a fair value of $709,000 on January 1, 2017, and an estimated economic life of 10 years. 6. Elmer incurs executory costs of $8,300 per year. (Use Accounts Payable) Prepare the journal entries for both the lessee and the lessor for 2017 to reflect the sale-leaseback agreement. No uncertainties exist, and collectibility is reasonably certain. To record amortization of profit on sale use Depreciation Expense account and not Sales Revenue account. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select No Entry for the account titles and enter o for the amounts. Round answers to o decimal places, e.g. 50,250.) Click here to view factor tables Debit Credit Date Account Titles and Explanation Elmers Restaurants (Lessee) (To record sale.) To record the lease.)
Answer #1

Answer:

journal entries in the books of Elmer's restaurant (lessee) :-
Date account tile & explanation debit($) credit($)
01-01-2017 cash               7,09,000
      Equipment     6,36,000
      Unearned profit on sale-lease        73,000
(To record sale)
01-01-2017 leased equipment (note 1) $6,51,954
         Lease liability (note 1) $6,51,954
(To record the lease)
Throughout 2017 Executory cost                    8,300
        Accounts payable          8,300
(To record executory cost )
12/31/2017 unearned profit on sale (note 4)                    7,300
          Depreciation expense          7,300
(To record amortization of profit on sale )
12/31/2017 depreciation exp (note 3)                  65,195
         Accumulated depreciation        65,195
(To record depreciation exp)
12/31/2017 Interest exp (see note 2)                  85,080
Lease liability (see note 2)                  30,306
         Cash     1,15,386
(To record payment )
Note :-
1 . Lease liability will be present value of total lease payments discounted at the borrowing rate of Elmer . Which will be as follows :-
115,386 x pvaf (12%,10)
=115,386*5.6502
Lease liability = $6,51,954.
2 . Intrest expense and lease liability to be debited is calculated as :-
Interest exp will be calculated on amount outstanding i.e. 709,000
Interest exp = 709,000x12% = $85,080
Lease liability to be debited = lease payment for the year - interest exp for the year
= 115,386-85,080 = $30,306
3 . Depreciation exp = 6,51,954/10 = $65,195
4 . Profit is to be amortized in the ratios of depreciation i.e. slm basis . Hence profit to be amortized = 73,000/10 = $7,300
2 . Journal entries in the books of liquidity finance co (lessor)
Date account tile & explanation debit($) Credit($)
01-01-2017 equipment               7,09,000
           Cash     7,09,000
(To record purchase)
Lease receivable               7,09,000
            Equipment     7,09,000
(To record the leaseback)
12/31/17 cash               1,15,386
           Lease receivable (115,386-70,900)        44,486
            Interest revenue        70,900
Note :-
1 . Lease receivable amounting = present value of lease payment discounted @ required rate of return
   = 115,386 x pvaf(10%,10)
= 115,386x6.1446 = 709,000
2. Interest = 709,000x10% = 70,900